# Random expansions of trees with bounded height

### v1

@@ 3. Probability logic with aggregation functions
We consider a logic that we call probability logic with aggregation functions, or $PLA^{*}$, where $PLA^{*}$ is a more general version of $PLA$ and $PLA^{+}$ which where considered in [21] and [22], respectively, and of the probability logic considered by Jaeger in [14, 15]. $PLA^{*}$ is a logic with (truth) values in the unit interval $[0,1]$. With $PLA^{*}$ we can, for example, express the proportion of all elements in a domain that have some property, and the proportion need not be 0 or 1. As a nontrivial example of what $PLA^{*}$ can express, Example 3.12 shows that the PageRank can be expressed with $PLA^{*}$. On finite structures, all queries that can be expressed by first-order logic can also be expressed by $PLA^{*}$ (as stated by Lemma 3.13).
@@
We exemplify what can be expressed with $PLA^{*}(\sigma)$, provided that it contains a binary relation symbol, with the notion of PageRank [6]. The PageRank of an internet site can be approximated in “stages” as follows (if we supress the “damping factor” for simplicity), where $IN_{x}$ is the set of sites that link to $x$, and $OUT_{y}$ is the set of sites that $y$ link to:
@@
The next basic lemma has analogue in first-order logic and is proved straightforwardly by induction on the complexity of $PLA^{*}$-formulas.
@@ 4. A general method for asymptotic elimination of aggregation functions
Let $\sigma$ be a finite and relational signature. In all of this section let $D_{n}$, $n\in\mathbb{N}^{+}$, be finite sets such that $\lim_{n\to\infty}|D_{n}|=\infty$, and let $\mathbf{W}_{n}$ be a set of $\sigma$-structures with domain $D_{n}$. We will describe a method for “asymptotically eliminating aggregation functions” if
  • the aggregation functions in question are “sufficiently” continuous, and
  • if there are 0/1-valued formulas with certain properties, including that the proportions of some definable sets “stabilize” as the domain size tends to infinity.
This method is studied in more detail by Koponen and Weitkämper in [23] and the definitions below and Theorem 4.8 come from [23]. Theorem 4.8 will be used later (in Section 7.5) to prove our main results.
@@
$\lim_{n\to\infty}\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\text{ for all $\bar{a}\in(B_{n})^{|\bar{x}|}$, $|\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}(\psi(\bar{a}))|\leq\varepsilon$}\big{\}}\Big{)}\ =\ 1.$
@@
To define the notions of continuity that we will use we need the notion of convergence testing sequence. It generalizes a similar notion used by Jaeger in [14]. The intuition is that a sequence $\bar{p}_{n}\in[0,1]^{<\omega}$, $n\in\mathbb{N}^{+}$, is convergence testing for parameters $c_{1},\ldots,c_{k}\in[0,1]$ and $\alpha_{1},\ldots\alpha_{k}\in[0,1]$ if the length of $\bar{p}_{n}$ tends to infinity as $n\to\infty$ and, as $n\to\infty$, all entries in $\bar{p}_{n}$ congregate ever closer to the “convergence points” in $\{c_{1},\ldots,c_{k}\}$, and the proportion of entries in $\bar{p}$ which congregate to $c_{i}$ converges to $\alpha_{i}$.
@@
(ii) We call $F$ admissible if $F$ is ct-continuous with respect to every choice of parameters $c_{i,j},\alpha_{i,j}\in[0,1]$, $i=1,\ldots,m$ and $j=1,\ldots,k_{i}$ (for arbitrary $m$ and $k_{i}$) such that $\alpha_{i,j}>0$ for all $i$ and $j$.
@@
The method that we consider for asymptotically eliminating continuous or admissible aggregation functions can be applied if one can find sets $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ of 0/1-valued formulas that satisfy the conditions of the following assumption:
@@
(ii) Suppose, in addition, that if $\varphi_{j}(\bar{x},\bar{y})$, $\chi_{j}(\bar{x},\bar{y})$, $\varphi^{\prime}_{i}(\bar{x})$, $\mathbf{Y}^{\varepsilon}_{n}$ and $\alpha_{i,j}$ are like in part 2 of Asumption 4.7 and $\alpha_{i,j}=0$, then, for all sufficiently large $n$, all $\bar{a}\in(D_{n})^{|\bar{x}|}$ and all $\mathcal{A}\in\mathbf{Y}^{\varepsilon}_{n}$, if $\mathcal{A}\models\varphi^{\prime}_{i}(\bar{a})$ then $\varphi_{j}(\bar{a},\mathcal{A})\cap\chi_{j}(\bar{a},\mathcal{A})=\emptyset$. If $F$ is admissible then $\varphi(\bar{x})$ is asymptotically equivalent to an $L_{0}$-basic formula with respect to $\mathbb{P}$.
@@ 5. Directed acyclic graphs, trees and closure types
For the rest of this article we let $\tau=\{E\}$ where $E$ is a binary relation symbol. Moreover, $\sigma$ will always denote a finite relational signature such that $\tau\subseteq\sigma$, although this later assumption will be repeated.
@@
(iii) Let $a,b\in G$. A directed path (of length $l$) from $a$ to $b$ is a sequence $c_{0},c_{1},\ldots,c_{l}\in T$ such that $l\geq 1$, $\mathcal{T}\models E(c_{i},c_{i+1})$ for all $i=0,\ldots,l-1$, $a=c_{0}$ and $b=c_{l}$. If there is a directed path from $a$ to $b$ then $a$ is called an ancestor of $b$ and $b$ is called a successor of $a$.
@@
    there is a unique element $a\in T$, called the root, such that $\mathcal{T}\models\neg\exists xE(x,a)$, and
@@
    for all $a\in T$, if $a$ is not the root then there is a unique $b\in T$, called the parent of $a$, such that $\mathcal{T}\models E(b,a)$, and in this case $a$ is called a child of $b$.
@@
Suppose that $\mathcal{T}$ is a tree, $a\in T$, $\{b_{1},\ldots,b_{k}\}\subseteq T$ and $a\in\mathrm{cl}_{\mathcal{T}}(b_{1},\ldots,b_{k})$. Then either $a=b_{i}$ for some $i$, or for some $l\in\mathbb{N}^{+}$ and $i$ there is a directed path from $a$ to some $b_{i}$ of length $l$. The statement “there is a directed path from $x$ to $y$ of length $l$” can be expressed by a first-order formula, say $\xi_{l}(x,y)$. We will use the expression ‘$x\in\mathrm{cl}(y)$’ to denote a formula $\xi_{l}(x,y)$ for some unspecified $l$, and more generally, the expression ‘$x\in\mathrm{cl}(y_{1},\ldots,y_{k})$’ means $\bigvee_{i=1}^{k}x\in\mathrm{cl}(y_{i})$.
@@
If, moreover, for every $R\in\sigma$, or arity $r$ say, and all $i_{1},\ldots,i_{r}\in\{1,\ldots,k\}$, either $R(x_{i_{1}},\ldots,x_{i_{r}})$ or $\neg R(x_{i_{1}},\ldots,x_{i_{r}})$ is a conjunct of $\varphi(x_{1},\ldots,x_{k})$, then we call $\varphi(x_{1},\ldots,x_{k})$ a complete atomic type over $\sigma$ with respect to $\mathcal{A}$. If the particular tree $\mathcal{T}$ is not important in the context we may omit the phrase ‘with respect to $\mathcal{T}$’.
@@
(i) The restriction of $p(\bar{x})$ to $\sigma^{\prime}$, denoted $p\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$, is a closure-type $p^{\prime}(\bar{x})$ over $\sigma^{\prime}$ (with respect to $\mathcal{T}$) such that $p(\bar{x})\models p^{\prime}(\bar{x})$, and for every closure-type $p^{*}(\bar{x})$ over $\sigma^{\prime}$, if $p(\bar{x})\models p^{*}(\bar{x})$ then $p^{\prime}(\bar{x})\models p^{*}(\bar{x})$.
(ii) The restriction of $p(\bar{x})$ to $\bar{y}$, denoted $p\negthickspace\upharpoonright\negthickspace\bar{y}$, is a closure type over $\sigma$ (with respect to $\mathcal{T}$) $p^{\prime}(\bar{y})$ such that $p(\bar{x})\models p^{\prime}(\bar{y})$, and for every closure-type $p^{*}(\bar{y})$ over $\sigma$, if $p(\bar{x})\models p^{*}(\bar{y})$ then $p^{\prime}(\bar{y})\models p^{*}(\bar{y})$.
It is not hard to see that, under the same assumptions as in the above definition, if $p(\bar{x})$ is a closure type over $\sigma$ with respect to the tree $\mathcal{T}$, then $p\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$ is a closure-type over $\sigma^{\prime}$ with respect to $\mathcal{T}$, and $p\negthickspace\upharpoonright\negthickspace\bar{y}$ are closure type over $\sigma$ with respect to $\mathcal{T}$. It will be convenient to use ‘$\models_{tree}$’ to denote consequence restricted to expansions of trees, as defined below.
@@
Suppose that $\tau\subseteq\sigma$ and let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$ are $0/1$-valued formulas. By the expression $\varphi(\bar{x})\models_{tree}\psi(\bar{x})$ we mean that if $\mathcal{A}$ is a $\sigma$-structure that expands a tree, $\bar{a}\in(A)^{|\bar{x}|}$, and $\mathcal{A}\models\varphi(\bar{a})$, then $\mathcal{A}\models\psi(\bar{a})$. The expression $\models_{tree}\psi(\bar{x})$ means the same as $\top\models_{tree}\psi(\bar{x})$, or equivalently, that $\forall\bar{x}\psi(\bar{x})$ is true in all $\sigma$-structures that expand a tree.
@@
Then $p^{*}(\bar{x},\bar{y})$ is a closure type over $\sigma$ with empty sequence of quantifiers in the beginning. Since $p^{*}(\bar{x},\bar{y})$ implies that ‘$\bar{x}\bar{y}$ is closed’ it follows that $p^{*}$ is self-contained. It is evident that if $\mathcal{A}$ is a $\sigma$-structure then $\mathcal{A}\models\forall\bar{x}\big{(}p(\bar{x})\leftrightarrow\exists\bar{y}p^{*}(\bar{x},\bar{y})\big{)}$. Suppose, moreover, that $\mathcal{A}$ expands a tree and $\mathcal{A}\models p^{*}(\bar{a},\bar{b})$. Then, for every $b\in\mathrm{rng}(\bar{b})$, either $b\in\mathrm{rng}(\bar{a})$ or there is a directed path of a particular length, say $l_{b}$, from $b$ to some $a_{b}\in\mathrm{rng}(\bar{a})$. In the later case, since $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ is a tree, it follows that $b$ is the unique element such that there is a directed path from $b$ to $a_{b}$ of length $l_{b}$. Therefore $\bar{b}$ is the unique tuple such that $\mathcal{A}\models p^{*}(\bar{a},\bar{b})$. $\square$
@@
Suppose that $\tau\subseteq\sigma$ and $p(\bar{x})$ is a closure-type over $\sigma$ where $\bar{x}=(x_{1},\ldots,k)$.
@@
In the next lemma, and later, the expression ‘$\mathrm{rng}(\bar{y})\subseteq\mathrm{cl}(\bar{x})$’ means that ‘for all $y_{i}\in\mathrm{rng}(\bar{y})$, $y_{i}\in\mathrm{cl}(\bar{x})$’.
@@
Let $\tau\subseteq\sigma$. A closure-basic formula over $\sigma$ is a formula of the form
$\bigwedge_{i=1}^{k}(\varphi_{i}(\bar{x})\rightarrow c_{i})$ where for each $i=1,\ldots,k$, $\varphi_{i}(\bar{x})$ is a complete closure type over $\sigma$ and $c_{i}\in[0,1]$.
@@
Proof. (i) Suppose that $\varphi_{i}(\bar{x})$, $i=1,\ldots,k$, is a closure-basic formula over $\sigma$. Let $q_{1}(\bar{x}),\ldots,q_{m}(\bar{x})$ enumerate, up to logical equivalence, all complete closure-basic types over $\sigma$ in the free variables $\bar{x}$. Suppose that $\mathcal{A}$ is a finite $\sigma$-structure abd $\bar{a}\in A^{|\bar{x}|}$. Observe that for each $i$ the value $\mathcal{A}(\varphi_{i}(\bar{a}))$ depends only on which $q_{j}(\bar{x})$ the sequence $\bar{a}$ satisfies. So let $c_{i,j}=\mathcal{A}(\varphi_{i}(\bar{a}))$ if $\mathcal{A}\models q_{j}(\bar{a})$. Then let $d_{j}=\mathsf{C}(c_{1,j},\ldots,c_{k,j})$ for $j=1,\ldots,m$. Now $\varphi(\bar{x})$ is equivalent in to the closure-basic formula $\bigwedge_{j=1}^{m}(q_{j}(\bar{x})\to d_{j})$.

### v2

@@ 3. Probability logic with aggregation functions
We consider a logic that we call probability logic with aggregation functions, or $PLA^{*}$, where $PLA^{*}$ is a more general version of $PLA$ and $PLA^{+}$ which were considered in [21] and [22], respectively, and of the probability logic considered by Jaeger in [14, 15]. $PLA^{*}$ is a logic with (truth) values in the unit interval $[0,1]$. With $PLA^{*}$ we can, for example, express the proportion of all elements in a domain that have some property, and the proportion need not be 0 or 1. As a nontrivial example of what $PLA^{*}$ can express, Example 3.12 shows that the PageRank can be expressed with $PLA^{*}$. On finite structures, all queries that can be expressed by first-order logic can also be expressed by $PLA^{*}$ (as stated by Lemma 3.13).
@@
We exemplify what can be expressed with $PLA^{*}(\sigma)$, provided that it contains a binary relation symbol, with the notion of PageRank [6]. The PageRank of an internet site can be approximated in “stages” as follows (if we supress the “damping factor” for simplicity), where $IN_{x}$ is the set of sites that link to $x$, and $OUT_{y}$ is the set of sites that $y$ links to:
@@
The next basic lemma has an analog in first-order logic and is proved straightforwardly by induction on the complexity of $PLA^{*}$-formulas.
@@ 4. A general method for asymptotic elimination of aggregation functions
Let $\sigma$ be a finite and relational signature. In all of this section let $D_{n}$, $n\in\mathbb{N}^{+}$, be finite sets such that $\lim_{n\to\infty}|D_{n}|=\infty$, and let $\mathbf{W}_{n}$ be a set of $\sigma$-structures with domain $D_{n}$. We will describe a method for “asymptotically eliminating aggregation functions” which is studied in more detail by Koponen and Weitkämper in [23]. The definitions below and Theorem 4.8 come from [23] where Theorem 4.8 will be used later (in Section 7.5) to prove our main results.
@@
$\lim_{n\to\infty}\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\text{ for all $\bar{a}\in(D_{n})^{|\bar{x}|}$, $|\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}(\psi(\bar{a}))|\leq\varepsilon$}\big{\}}\Big{)}\ =\ 1.$
@@
To define the notions of continuity that we will use we need the notion of convergence testing sequence which generalizes a similar notion used by Jaeger in [14].
@@
(ii) We call $F$ admissible if $F$ is ct-continuous with respect to every choice of parameters $c_{i,j},\alpha_{i,j}\in[0,1]$, $i=1,\ldots,m$ and $j=1,\ldots,k_{i}$ (for arbitrary $k_{i}$) such that $\alpha_{i,j}>0$ for all $i$ and $j$.
@@
The aggregation functions am, gm and $\mathrm{length}^{-\beta}$ are continuous, which is proved in [21] in the case of am and gm. In the case of $\mathrm{length}^{-\beta}$ the claim is easy to prove. The aggregation functions max and min are admissible (which is proved in [21]) but not continuous. To see that max is not continuous, consider, for $n\in\mathbb{N}$, $\bar{p}_{n}$ and $\bar{q}_{n}$, both of length $n+1$, where all entries of $\bar{p}_{n}$ are 0, the first entry of $\bar{q}_{n}$ is 1 and the rest of the entries are 0. Let $c_{1}=0$, $c_{2}=1$, $\alpha_{1}=1$, and $\alpha_{2}=0$. It is straightforward to verify that both $(\bar{p}_{n}:n\in\mathbb{N})$ and $(\bar{q}_{n}:n\in\mathbb{N})$ are convergence testing with parameters $c_{1},c_{2}$ and $\alpha_{1},\alpha_{2}$. But clearly $\max(\bar{p}_{n})=0$ and $\max(\bar{q}_{n})=1$ for all $n$, so $|\max(\bar{p}_{n})-\max(\bar{q}_{n})|$ does not tend to 0 as $n\to\infty$.
@@
The method that we consider for asymptotically eliminating continuous or admissible aggregation functions can be applied if one can find sets $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ of 0/1-valued formulas that satisfy the conditions of Assumption 4.7 below (which comes from [23]). The intuition behind the technical part (2) of the assumption is that for the sets $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ and every $\varphi(\bar{x},\bar{y})\in L_{0}$ there is a set $L_{\varphi(\bar{x},\bar{y})}\subseteq L_{1}$ of formulas defining some “allowed” conditions (with respect to $\varphi(\bar{x},\bar{y})$) and there are some $\varphi^{\prime}_{1}(\bar{x}),\ldots,\varphi^{\prime}_{s}(\bar{x})\in L_{0}$ such that if $\mathcal{A}\models\varphi^{\prime}_{i}(\bar{a})$ and $\chi(\bar{x},\bar{y})\in L_{\varphi(\bar{x},\bar{y})}$, then the fraction $|\varphi(\bar{a},\mathcal{A})\cap\chi(\bar{a},\mathcal{A})|/|\chi(\bar{a},\mathcal{A})|$ is with high probability close to a number $\alpha_{i}$ that depends only on $\varphi(\bar{x},\bar{y}),\chi(\bar{x},\bar{y}),\varphi_{i}(\bar{x})$ and the sequence of probability distributions $\mathbb{P}$. As we allow aggregation functions with arity $m>1$, part (2) needs to simultaneously speak of a sequence $\varphi_{1}(\bar{x},\bar{y}),\ldots,\varphi_{m}(\bar{x},\bar{y})\in L_{0}$.
@@
(ii) Suppose, in addition, that the following holds if $\varphi_{j}(\bar{x},\bar{y})\in L_{0}$, $\chi_{j}(\bar{x},\bar{y})\in L_{1}$, $\varphi^{\prime}_{i}(\bar{x})\in L_{0}$, $\mathbf{Y}^{\varepsilon}_{n}$ and $\alpha_{i,j}$ are like in part 2 of Asumption 4.7: If $\alpha_{i,j}=0$ then, for all sufficiently large $n$, all $\bar{a}\in(D_{n})^{|\bar{x}|}$ and all $\mathcal{A}\in\mathbf{Y}^{\varepsilon}_{n}$, if $\mathcal{A}\models\varphi^{\prime}_{i}(\bar{a})$ then $\varphi_{j}(\bar{a},\mathcal{A})\cap\chi_{j}(\bar{a},\mathcal{A})=\emptyset$. Then it follows that if $F$ is admissible then $\varphi(\bar{x})$ is asymptotically equivalent to an $L_{0}$-basic formula with respect to $\mathbb{P}$.
@@ 5. Directed acyclic graphs, trees and closure types
For the rest of this article we let $\tau=\{E\}$ where $E$ is a binary relation symbol. Moreover, $\sigma$ will always denote a finite relational signature such that $\tau\subseteq\sigma$.
@@
(iii) Let $a,b\in G$. A directed path (of length $l$) from $a$ to $b$ is a sequence $c_{0},c_{1},\ldots,c_{l}\in T$ such that $l\geq 1$, $\mathcal{G}\models E(c_{i},c_{i+1})$ for all $i=0,\ldots,l-1$, $a=c_{0}$ and $b=c_{l}$. If there is a directed path from $a$ to $b$ then $a$ is called an ancestor of $b$ and $b$ is called a successor of $a$.
@@
    there is a unique element $a\in T$, called the root, such that $\mathcal{T}\models\neg\exists xE(x,a)$,
@@
    for all $a\in T$, if $a$ is not the root then there is a unique $b\in T$, called the parent of $a$, such that $\mathcal{T}\models E(b,a)$, and in this case $a$ is called a child of $b$, and
  4. (4)
    there do not exist $n\in\mathbb{N}^{+}$ and distinct $a_{0},\ldots,a_{n}\in T$ such that $\mathcal{T}\models E(a_{n},a_{0})$ and $\mathcal{T}\models E(a_{k},a_{k+1})$ for all $k=0,\ldots,n-1$.
@@
Suppose that $\mathcal{T}$ is a tree, $a\in T$, $\{b_{1},\ldots,b_{k}\}\subseteq T$ and $a\in\mathrm{cl}_{\mathcal{T}}(b_{1},\ldots,b_{k})$. Then either $a=b_{i}$ for some $i$, or for some $l\in\mathbb{N}^{+}$ and $i$ there is a directed path from $a$ to some $b_{i}$ of length $l$. The statement “there is a directed path from $x$ to $y$ of length at most $l$” can be expressed by a first-order formula, say $\xi_{l}(x,y)$. It follows that in all trees with height at most $l$ the property ‘$x\in\mathrm{cl}(y)$’ is expressed by the formula $\xi_{l}(x,y)$.
@@
For the rest of this section we fix some $\Delta\in\mathbb{N}^{+}$ and we assume that all trees mentioned have height at most $\Delta$, so “tree” will mean “tree of height at most $\Delta$”. With this assumption the property ‘$x\in\mathrm{cl}(y)$’ is expressed by the formula $\xi_{\Delta}(x,y)$ from Remark 5.6 and we will use the more intuitive expression ‘$x\in\mathrm{cl}(y)$’ to denote that formula. More generally, the expression ‘$x\in\mathrm{cl}(y_{1},\ldots,y_{k})$’ will mean ‘$\bigvee_{i=1}^{k}x\in\mathrm{cl}(y_{i})$’.
@@
If, moreover, for every $R\in\sigma$, of arity $r$ say, and all $i_{1},\ldots,i_{r}\in\{1,\ldots,k\}$, either $R(x_{i_{1}},\ldots,x_{i_{r}})$ or $\neg R(x_{i_{1}},\ldots,x_{i_{r}})$ is a conjunct of $\varphi(x_{1},\ldots,x_{k})$, then we call $\varphi(x_{1},\ldots,x_{k})$ a complete atomic type over $\sigma$ with respect to $\mathcal{A}$. If the particular tree $\mathcal{T}$ is not important in the context we may omit the phrase ‘with respect to $\mathcal{T}$’.
@@
(i) The restriction of $p(\bar{x})$ to $\sigma^{\prime}$, denoted $p\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$, is a closure-type $p^{\prime}(\bar{x})$ over $\sigma^{\prime}$ (with respect to $\mathcal{T}$) such that $p(\bar{x})\models p^{\prime}(\bar{x})$, and for every closure-type $p^{*}(\bar{x})$ over $\sigma^{\prime}$, if $p(\bar{x})\models p^{*}(\bar{x})$ then $p^{\prime}(\bar{x})\models p^{*}(\bar{x})$. (The restriction is not syntactically unique, but it is unique up to logical equivalence, so technically we just choose one of the formulas that satisfy the condition of being a restriction.)
(ii) The restriction of $p(\bar{x})$ to $\bar{y}$, denoted $p\negthickspace\upharpoonright\negthickspace\bar{y}$, is a closure type over $\sigma$ (with respect to $\mathcal{T}$) $p^{\prime}(\bar{y})$ such that $p(\bar{x})\models p^{\prime}(\bar{y})$, and for every closure-type $p^{*}(\bar{y})$ over $\sigma$, if $p(\bar{x})\models p^{*}(\bar{y})$ then $p^{\prime}(\bar{y})\models p^{*}(\bar{y})$. (Again, the restriction is unique up to logical equivalence.)
It is not hard to see that, under the same assumptions as in the above definition, if $p(\bar{x})$ is a closure type over $\sigma$ with respect to the tree $\mathcal{T}$, then $p\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$ is a closure-type over $\sigma^{\prime}$ with respect to $\mathcal{T}$, and $p\negthickspace\upharpoonright\negthickspace\bar{y}$ is a closure type over $\sigma$ with respect to $\mathcal{T}$. It will be convenient to use ‘$\models_{tree}$’ to denote consequence restricted to expansions of trees, as defined below.
@@
Suppose that $\tau\subseteq\sigma$ and let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$ be $0/1$-valued formulas. By the expression $\varphi(\bar{x})\models_{tree}\psi(\bar{x})$ we mean that if $\mathcal{A}$ is a $\sigma$-structure that expands a tree (of height at most $\Delta$, $\bar{a}\in(A)^{|\bar{x}|}$, and $\mathcal{A}\models\varphi(\bar{a})$, then $\mathcal{A}\models\psi(\bar{a})$. The expression $\models_{tree}\psi(\bar{x})$ means the same as $\top\models_{tree}\psi(\bar{x})$, or equivalently, that $\forall\bar{x}\psi(\bar{x})$ is true in all $\sigma$-structures that expand a tree.
@@
Then $p^{*}(\bar{x},\bar{y})$ is a closure type over $\sigma$ with empty sequence of quantifiers in the beginning. Since $p^{*}(\bar{x},\bar{y})$ implies that ‘$\bar{x}\bar{y}$ is closed’ it follows that $p^{*}$ is self-contained. It is evident that if $\mathcal{A}$ is a $\sigma$-structure then $\mathcal{A}\models\forall\bar{x}\big{(}p(\bar{x})\leftrightarrow\exists\bar{y}p^{*}(\bar{x},\bar{y})\big{)}$. Suppose, moreover, that $\mathcal{A}$ expands a tree and $\mathcal{A}\models p^{*}(\bar{a},\bar{b})$. Then, for every $b\in\mathrm{rng}(\bar{b})$, either $b\in\mathrm{rng}(\bar{a})$ or there is a directed path of a particular length, say $l_{b}$, from $b$ to some $a_{b}\in\mathrm{rng}(\bar{a})$. In the latter case, since $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ is a tree, it follows that $b$ is the unique element such that there is a directed path from $b$ to $a_{b}$ of length $l_{b}$. Therefore $\bar{b}$ is the unique tuple such that $\mathcal{A}\models p^{*}(\bar{a},\bar{b})$. $\square$
@@
Suppose that $\tau\subseteq\sigma$ and $p(\bar{x})$ is a closure-type over $\sigma$ where $\bar{x}=(x_{1},\ldots,x_{k})$.
@@
In the next lemma, and later, if $\bar{y}=(y_{1},\ldots,y_{k})$ the expression ‘$\mathrm{rng}(\bar{y})\subseteq\mathrm{cl}(\bar{x})$’ denotes the formula ‘$\bigwedge_{i=1}^{k}y_{i}\in\mathrm{cl}(\bar{x})$’.
@@
Let $\tau\subseteq\sigma$. A closure-basic formula over $\sigma$ is a formula of the form $\bigwedge_{i=1}^{k}(\varphi_{i}(\bar{x})\rightarrow c_{i})$ where for each $i=1,\ldots,k$, $\varphi_{i}(\bar{x})$ is a complete closure type over $\sigma$ and $c_{i}\in[0,1]$.
@@
Proof. (i) Suppose that $\varphi_{i}(\bar{x})$, $i=1,\ldots,k$, is a closure-basic formula over $\sigma$. Let $q_{1}(\bar{x}),\ldots,q_{m}(\bar{x})$ enumerate, up to logical equivalence, all complete closure types over $\sigma$ in the free variables $\bar{x}$. Suppose that $\mathcal{A}$ is a finite $\sigma$-structure and $\bar{a}\in A^{|\bar{x}|}$. Observe that for each $i$ the value $\mathcal{A}(\varphi_{i}(\bar{a}))$ depends only on which $q_{j}(\bar{x})$ the sequence $\bar{a}$ satisfies. So let $c_{i,j}=\mathcal{A}(\varphi_{i}(\bar{a}))$ if $\mathcal{A}\models q_{j}(\bar{a})$. Then let $d_{j}=\mathsf{C}(c_{1,j},\ldots,c_{k,j})$ for $j=1,\ldots,m$. Now $\varphi(\bar{x})$ is equivalent to the closure-basic formula $\bigwedge_{j=1}^{m}(q_{j}(\bar{x})\to d_{j})$.